A sample of N2 has 1.70 moles and occupies 3.80 l at 25.0˚C.
A) What volume will it occupy with 2.60 moles?
B) How many moles are in a sample that occupies 1.45 L?
1
Expert's answer
2013-03-05T08:54:45-0500
According to the ideal gas law:

pV=nRT

p – pressure (atm)

V – volume (L)

n – amount (moles)

R – universal gas constant

T – temperature (K)

P, R, T – don’t change, that’s why we can writea proportion

V1/V2 = n1/n2

a) What volume will it occupy with 2.60 moles?
V2 = V1 x n2 / n1
V2 = 3.80 x 2.60 / 1.70 = 5.81 L
Answer: V2 = 5.81 L

b) How many moles are in a sample that occupies 1.45 L?
n2 = n1 x V2 / V1 = 1.70 x 1.45 / 3.80 = 0.65 mol
Answer: n2 = 0.65 mol